the plane surface of a plano convex lens of focal length 20cm is silvered .it will behaves as

Grade:12

1 Answers

Balaji
6 years ago
1/FL= (n-1)[1/R-0] hear FL is focal length of lens Then power of lens Pl=1/FLFM=infinity hear FM is focal length of mirror Then power of mirror is zero Total power is P=pl+pm+plLens will behave as a concave mirror of focal length R/[2(n-1)]hear n is refractive index